# Time to Retrain? Detecting Concept Drifts in Machine Learning Systems

### v1

@@ Abstract
With the boom of machine learning (ML) techniques, software practitioners build ML systems to process the massive volume of streaming data for diverse software engineering tasks such as failure prediction in AIOps. Trained using historical data, such ML models encounter performance degradation caused by concept drift, i.e., data and inter-relationship (concept) changes between training and production. It is essential to use concept rift detection to monitor the deployed ML models and re-train the ML models when needed.
In this work, we explore applying state-of-the-art (SOTA) concept drift detection techniques on synthetic and real-world datasets in an industrial setting. Such an industrial setting requires minimal manual effort in labeling and maximal generality in ML model architecture. We find that current SOTA semi-supervised methods not only require significant labeling effort but also only work for certain types of ML models.
To overcome such limitations, we propose a novel model-agnostic technique (CDSeer) for detecting concept drift. Our evaluation shows that CDSeer has better precision and recall compared to the state-of-the-art while requiring significantly less manual labeling. We demonstrate the effectiveness of CDSeer at concept drift detection by evaluating it on eight datasets from different domains and use cases. Results from internal deployment of CDSeer on an industrial proprietary dataset show a 57.1% improvement in precision while using 99% fewer labels compared to the SOTA concept drift detection method. The performance is also comparable to the supervised concept drift detection method, which requires 100% of the data to be labeled. The improved performance and ease of adoption of CDSeer are valuable in making ML systems more reliable.
@@ I Introduction
Modern systems generate huge amounts of data and are expected to reach 463 exabytes by 2025 [1] where a significant portion of the data comes in the form of time series data streams. The data can come from sensor recordings and signals from internet-of-things (IoT) and autonomous driving systems (ADS) [1]; or alerts, logs, and measurements from software systems [2, 3, 4, 5]. As these systems automatically produce data at high volume and velocity, posing a significant challenge to manual processing capabilities, practitioners develop machine learning (ML) systems powered by ML models to perform critical tasks such as predicting software failures [6].
@@
One open challenge to the reliability of deployed ML systems is that the ML models inevitably become outdated and experience degraded performance on new data due to the constantly changing nature of real-world environments and data distributions. Such differences can be categorized as data or concept drift. Data drift occurs when production data follows a different distribution from training data or is unseen by the model during training. Conversely, concept drifts happen where the data features are the same but the relationship between the features and the class (concepts) changes. Hence, it is important to continuously test and monitor a deployed ML system for both data and concept drift. Once a drift is detected, maintainers of ML systems should be alerted that a deployed ML model is obsolete and it is time to retrain.
@@
Current research efforts have been made to detect data drifts and concept drift detection (CDD) [7, 8, 9]. Supervised CDD is impractical as manual labeling true labels on all data is time-consuming and costly, especially when a domain expert is needed. Unsupervised CDD [10] relies on detecting data shifts and therefore cannot detect CDD efficiently. Instead, semi-supervised concept drift detection is more desired and practical due to significantly reduced manual labeling effort [8, 9]. However, such effort reduction can be affected by the frequency and the amount of data a system receives. In real-world systems, data can arrive in large volumes and at unpredictable rates, making it challenging to identify drift accurately without overwhelming computational resources or manually labeling ground truth [8, 9]. Additionally, the varying quality and representativeness of incoming data can impact the effectiveness of semi-supervised methods, necessitating strategies to adapt dynamically to changing data streams.
@@
Particularly, we applied the SOTA semi-supervised CDD technique, i.e., ECHO [8] on various datasets and models, including one internally in Ericsson’s setting, one dataset from Google on predicting job failure [11], and four datasets from prior works. Our experiment shows that the SOTA technique requires a non-trivial amount of manual labels up to 100%, calling for better CDD techniques. When we discuss these results on real datasets and their implications if the SOTA CDDs are applied in an industrial setting, we find the limitations make the SOTA CDDs unsuitable for real-world deployment. Considering these results on real datasets and their implications for applying state-of-the-art (SOTA) concept drift detectors (CDDs) in an industrial setting, we find that the limitations make these SOTA CDDs unsuitable for real-world deployment.
@@
We summarize three key industry concerns the SOTA techniques fail to address.
@@
  • Excessive labeling effort. The SOTA semi-supervised CDDs require an average of 5% and in some cases, up to 100% of the time series to be labeled [12]. While this is a huge improvement compared to supervised CDDs, this is still challenging to achieve in production.
@@
    Industry Need. Experts should be required to label as few data points as possible, and the points labeled should be as useful as possible, i.e., minimal manual labeling to achieve maximal accuracy in detecting concept drift.
@@
  • Lack of flexibility on which data points require labeling. The SOTA CDDs request manual labels individually on demand based on metrics from datasets and models, i.e., cannot tell when and which manual labeling is required beforehand. This prevents the planning of labeling efforts in advance.
@@
    Industry Need. Experts prefer to label multiple data points each time at their convenience. Also, experts prefer to select as few points to label as possible, preferably points selected at random so it costs them the least labeling effort.
@@
  • Lack of generality for different ML models. The SOTA CDDs are optimized for specific model architectures and cannot be generalized for a different model architecture.
@@
    Industry Need. There is much need for black-box CDD since the ML models can be already trained, and it is difficult to convince users to adopt a new model architecture especially if the current one is optimized for their use case.
@@
To address the challenges above, we propose CDSeer, a novel semi-supervised CDD technique. CDSeer significantly reduces the workload for practitioners in manual labeling and is model- and distribution-agnostic. CDSeer works with ML models of different architectures and is independent of underlying data distribution. CDSeer’s innovations are as follows.
@@
    Model- and data distribution-agnostic predictable sampling for manual labeling. CDSeer utilizes a new sampling process that removes the dependence on the model architecture and data specifics. Hence CDSeer can detect concept drifts when the SOTA fails. This sampling process is also predictable in terms of the number of labels requested and the frequency of such requests, allowing for the planning of labeling efforts.
@@
    Semi-supervised learning that utilizes an inspector model to provide pseudo-GT labels for concept-drift detection. We propose a novel method to acquire pseudo labels from an inspector model to substitute true labels as the ground truth labels for concept drift prediction, which significantly lowers the number of needed labels while maintaining competitive performance to SOTA.
@@
    Model-agnostic concept drift detection. Our proposed CDSeer can be deployed to monitor ML models regardless of their architectures. Comparing this to SOTA CDD methods [9, 8] which utilizes ensemble classifiers for both prediction and concept drift detection, which is not adaptable to pre-trained industry models.
Our contributions include (1) a novel semi-supervised CDD technique that outperforms the SOTA techniques and satisfies the restrictions in industry reality; and (2) a comprehensive evaluation of CDSeer. We evaluate CDSeer on publicly available synthetic/real datasets and a proprietary dataset internal to a telecommunications company. The study on the proprietary dataset shows that CDSeer leads to comparable performance improvement on model accuracy with the supervised PHT and outperforms the SOTA semi-supervised while using very few labeled data points.
Relevance to Software Engineering in Practice. First, serving ML systems and detecting performance degradation over time (such as model obsolete) are active research areas in the Software Engineering community (e.g., [6]). Our work focuses on monitoring/testing deployed ML systems from a practical perspective, an important yet overlooked research area. Second, research on concept drift techniques can be applied to strengthen downstream SE research. For instance, ML models built to predict production system failures and performance degradation suffer from concept drifts in [6, 13]. Third, we propose a novel technique CDSeer to fill the gap between academia and software industry reality. CDSeer is shown to outperform the SOTA and is deployed in the industry. Our work is well suited to the SEIP track.
@@ II-A Concept Drift and Data Drift
Readers can observe the mentioned variation in terminology in citations [23, 7, 24, 10]. Many methods claim concept drift detection but only monitor $P_{training}(X)\neq P_{testing}(X)$ under the assumption that concept drifts happen simultaneously with data drifts [16]. Hence, these methods cannot detect concept drift without data drifts. Our work focuses on detecting real concept drift when $P_{training}(y|X)\neq P_{testing}(y|X)$ and $P_{training}(X)=P_{testing}(X)$. We do not monitor data drift.
@@ II-B Semi-Supervised Concept Drift Detection for Machine Learning Systems
CDDs can be categorized into supervised, semi-supervised, and unsupervised. Supervised approaches require true labels for all data, which is unrealistic. Differently, unsupervised and semi-supervised approaches require true labels for zero or a few data points, thus more practical in real-world settings. The current unsupervised techniques target cases where concept and data drifts happen simultaneously, hence ineffective at detecting concept drift alone.
Based on our industry needs, we believe semi-supervised approaches are suitable to detect concept drift accurately and efficiently in production. Semi-supervised approaches require limited manual labeling and can target concept drifts alone.
In a nutshell, our work focuses on proposing a better semi-supervised approach to detect concept drifts as we believe semi-supervised approaches are better aligned with our industry needs.
@@ III-A Concept Drift Detection
Supervised concept drift detection. Supervised CDD methods use the true labels of all data points to detect concept drift. They either compare statistics of the classifier error rate between the current and reference windows [25, 26, 27, 28, 29, 30] or use the intersection of confidence intervals from multiple CDDs to detect concept drift [31, 32]. While these methods are fast and accurate, they require labels for all data points which is unrealistic in production given the velocity of data streams [33].
(Semi)-supervised concept drift detection. As a more production-friendly alternative, semi-supervised CDD methods solve the problem by requiring only a subset of the data to be labeled. Existing methods can do this by using active learning [34, 35, 36], computational geometry [37], class imbalance [38], classifier confidence [39, 8], pseudo-error from a classifier ensemble [9], grid density sampling [40], ambiguity concerning the decision boundary and neighboring points [12], etc. to select only important data instances for labeling, reducing the labeling effort. However, the reliance on specific metrics to sample labels is vulnerable to data distribution and model architecture changes. Also, requiring the original model to follow certain ML architectures significantly limits adoption. We look to overcome these limitations.
Unsupervised concept drift detection. Unsupervised CDDs methods [10] detect drifts offline in batch [41, 42, 43, 44, 45, 46, 47, 48, 49, 50] or online by instances [51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. These methods detect concept drift by monitoring $P_{training}(X)\neq P_{testing}(X)$ under the assumption that changes in the data distribution and the posterior probability distributions often happen simultaneously [16]. Hence, unsupervised methods cannot detect concept drift when data drift is absent (Subsection II-A).  CDSeer aims to detect real concept drift even when the input data distribution stays the same.
@@ III-B Testing Machine Learning Systems
Artificial Intelligence for IT Operations (AIOps) AIOps leverage ML models for tasks such as anomaly detection and failure prediction [62, 63, 64, 3, 65, 66, 67, 5, 2, 68] using logs, measurements, and spatial information. These models are highly susceptible to data evolution with impairs their performance and stability. Current mitigation strategies are retraining the models periodically [67, 68] or retraining the models upon alarms by supervised concept drift detectors [6]. Existing methods are very costly in manual efforts due to the amount of labeling needed. In this work, we explore CDSeer’s CDD for AIOps to reduce cost.
Model Robustness and Relevance Testing. Model testing detects mismatches between the model and data [69, 70] and ensures trained models’ resistance to data noises [71, 72, 73]. However, these techniques test pre-trained models offline and more online techniques are needed [61]. CDSeer detect concept drift online to test the robustness and relevance of ML models. Thus  CDSeer can benefit the utilization and deployment of ML models for SE tasks. In this work, we applied CDSeer to an internal dataset for anomaly detection in router systems and the Google Cluster Trace dataset for job failure prediction.
@@ IV CDSeer: Learning-based Concept Drift Detection
In this section, we first summarize our approach at a high level (Section IV-A). Then, we describe the framework step-by-step, with design choices and implementation details.
@@ IV-A An Overview of CDSeer
At a high level, CDSeer works as follows. As new data points arrive, CDSeer detects concept drift via an inspector model which predicts the (pseudo-) labels of the streaming data points in parallel with the online model prediction of labels. The distribution of the classification error between the online model’s predicted labels and the pseudo labels is monitored by a supervised concept drift detection module to detect the concept drifts. The window of the latest data points is then continuously clustered. Clustering allows the separation of the data points into strata where stratified samples are generated. Experts would then label these sampled points. Afterward, CDSeer stores these true labels in memory. Via semi-supervised learning, true labels in the memory are spread to the window of the newest data points, where these newly generated labels and their features are used to update the inspector model. This entire process is done online. Finally, while this is not the main focus of our experiments, we follow convention and retrain the online model when concept drift is detected using true labeled data.
@@ 16             InspectorModel $\leftarrow$ trainModel$(W,TrueLabels)$;
We utilize a CDD to monitor the distribution of the $error$s rate and alert when it crosses a threshold, signaling the online and inspector models are trained using different concepts. For this purpose, the CDD can be chosen from well-known CDDs such as the Drift Detection Method (DDM), PageHinkley Test (PHT), etc. The CDD method monitors the aggregated $error$ rate between the online and inspector models for increases surpassing a predefined threshold, which triggers a concept drift alarm between the online and the inspector models. Then, the framework records the corresponding instance and proceeds to the next data point.

### v2

@@ Abstract
With the boom of machine learning (ML) techniques, software practitioners are building ML systems to process massive volumes of streaming data for diverse software engineering tasks, such as failure prediction in AIOps. Trained on historical data, these models can suffer performance degradation due to concept drift, where the inter-relationship between features and labels (concepts) changes between training and production. Therefore, concept drift detection is essential to monitor deployed ML models and retrain them as needed.
In this work, we explore applying state-of-the-art (SOTA) semi-supervised concept drift detection techniques on synthetic and real-world datasets in an industrial setting, which requires minimal manual labeling and maximal model compatibility. We find that current SOTA methods not only require significant labeling effort but also are model-specific. To address these limitations, we propose CDSeer, a novel model-agnostic technique to detect concept drift.
Our evaluation shows that CDSeer outperforms the SOTA in precision and recall while requiring significantly less manual labeling. We demonstrate CDSeer’s effectiveness at concept drift detection by evaluating it on eight datasets from different domains and use cases. Internal deployment of CDSeer on a proprietary industrial dataset shows a 57.1% improvement in precision while using 99% fewer labels compared to the SOTA method. Its performance is also comparable to the supervised (fully labeled) concept drift detection method. The improved performance and ease of adoption make CDSeer valuable in enhancing the reliability of ML systems.
@@ I Introduction
Modern systems generate huge amounts of data and are expected to reach 463 exabytes by 2025 [1] where a significant portion of the data consists of time series data streams. The data can come from sensor recordings and signals from internet-of-things (IoT) and autonomous driving systems (ADS) [1]; or alerts, logs, and measurements from software systems [2, 3, 4, 5]. As these systems automatically produce data at high volume and velocity, posing a significant challenge to manual processing capabilities, practitioners develop machine learning (ML) systems powered by ML models to perform critical tasks such as predicting software failures [6].
@@
One open challenge to the reliability of deployed ML systems is that the ML models inevitably become outdated and experience degraded performance on new data due to the constantly changing nature of real-world environments and data distributions. Such differences can be categorized as data or concept drift. Data drift occurs when production data follows a different distribution from training data or is unseen by the model during training. Conversely, concept drift happens where the data features remain the same, but the relationship between the features and the class (concepts) changes. Hence, it is important to continuously test and monitor a deployed ML system for both data and concept drift. Once a drift is detected, maintainers of the ML system should be alerted that the deployed model is obsolete and it is time to retrain.
@@
Current concept drift detectors (CDDs) are classified into three types: supervised CDDs, which only use labeled data for detection; unsupervised CDDs, which only use unlabeled data; and semi-supervised CDDs, which use a mix of unlabeled and labeled data [7, 8, 9]. Supervised CDDs are impractical as manual labeling true labels for all data is time-consuming and costly, especially when a domain expert is needed. Unsupervised CDDs [10] rely on detecting data drifts to identify concept drifts; therefore, it cannot reliably detect concept drift in the absence of data drift. Instead, semi-supervised CDDs are preferred for their practicality, as they significantly reduce manual labeling effort [8, 9]. However, this reduction in effort can be influenced by the frequency and volume of data a system receives. In real-world systems, data can arrive in large volumes and at unpredictable rates, making it challenging to accurately identify drift without overwhelming computational resources or manually labeling ground truth (GT) [8, 9]. Additionally, the varying quality and representativeness of incoming data can impact the effectiveness of semi-supervised methods, necessitating strategies to adapt dynamically to changing data streams.
@@
Particularly, we applied ECHO [8], the SOTA semi-supervised CDD, on various datasets and models, including one internally in Ericsson’s setting, one dataset from Google on predicting job failure [11], and four datasets from prior works. Our experiment shows that the SOTA technique requires a non-trivial amount of manual labels up to 100%, calling for better concept drift detection techniques. When we discuss these results on real datasets and their implications if the SOTA CDDs are applied in an industrial setting, we find the limitations make the SOTA CDDs unsuitable for real-world deployment. Considering these results in real datasets and their implications for applying state-of-the-art (SOTA) CDDs in an industrial setting, we find that the limitations make these SOTA CDDs unsuitable for real-world deployment.
@@
We summarize three key industry concerns that SOTA techniques fail to address.
@@
  • Excessive labeling effort. The SOTA semi-supervised CDDs require an average of 5%, and up to 100% of the time series to be labeled [12]. While this represents a significant improvement over supervised CDDs, it remains challenging to achieve in production.
@@
    Industry Need. Experts should be required to label as few data points as possible, and the labeled points should be as useful as possible, i.e., minimizing manual labeling while maximizing accuracy in detecting concept drift.
@@
  • Lack of flexibility in selecting which data points require labeling. The SOTA CDDs request manual labels of specific data points individually and on-demand based on dataset and model metrics, i.e., cannot tell when and which labels are needed beforehand. This limits the ability to plan labeling efforts.
@@
    Industry Need. Experts prefer to label multiple data points in each session at their convenience. Also, they prefer to select as few points to label as possible, preferably points that cost them the least labeling effort.
@@
  • Lack of generality for different ML models. SOTA CDDs are optimized for specific model architectures and cannot be generalized to different model architectures.
@@
    Industry Need. There is much need for black-box (model architecture-agnostic) CDD since ML models can already be trained, and it is difficult to convince users to adopt a new model, especially if the current one is optimized for their use case.
@@
To address the challenges above, we propose CDSeer, a novel semi-supervised CDD. CDSeer significantly reduces manual labeling workload and is both model- and distribution-agnostic. CDSeer works with ML models of various architectures and is independent of the underlying data distribution. CDSeer’s innovations are as follows.
@@
    Model- and data distribution-agnostic predictable sampling for manual labeling. CDSeer employs a novel sampling process that eliminates reliance on model and data specifics, enabling it to detect concept drift where SOTA methods may fail. This process is also predictable in terms of label requests and their frequency, allowing for better planning of labeling efforts.
@@
    Semi-supervised learning using an inspector model to provide pseudo-GT labels for concept drift detection. We propose a novel method that leverages an inspector model to generate pseudo labels to replace true labels as the GT for concept drift detection. This significantly reduces the number of required labels while maintaining competitive performance compared to SOTA.
@@
    Model-agnostic concept drift detection. CDSeer can monitor ML models regardless of their architecture. In contrast, SOTA CDDs [9, 8] rely on ensemble classifiers for both prediction and concept drift detection, making them unsuitable for pre-trained industry models.
Our contributions include (1) a novel semi-supervised CDD that outperforms the SOTA techniques and satisfies industry constraints, and (2) a comprehensive evaluation of CDSeer. We evaluate CDSeer on publicly available synthetic and real datasets, as well as a proprietary dataset internal to Ericsson. The study on the proprietary dataset shows that CDSeer achieves comparable performance improvements in model accuracy to the supervised method and outperforms the SOTA semi-supervised method while using significantly fewer labeled data points.
Relevance to Software Engineering in Practice. First, serving ML systems and detecting performance degradation over time (such as model obsolescence) are active research areas in the Software Engineering community (e.g., [6]). Our work focuses on monitoring and testing deployed ML systems from a practical perspective, an important yet overlooked research area. Second, research on concept drift techniques can be applied to strengthen downstream SE research. For instance, ML models built to predict production system failures and performance degradation suffer from concept drifts in [6, 13]. Third, we propose a novel technique, CDSeer, to fill the gap between academic research and software industry practices. CDSeer is shown to outperform the SOTA and is deployed in the industry. Our work is well suited for the SEIP track.
@@ II-A Concept Drift and Data Drift
Readers can observe the variation in terminology in citations [23, 7, 24, 10]. Many methods claim concept drift detection but only monitor $P_{training}(X)\neq P_{testing}(X)$ under the assumption that concept drifts happen simultaneously with data drifts [16]. Hence, these methods cannot detect concept drift without data drifts. Our work focuses on detecting real concept drift when $P_{training}(y|X)\neq P_{testing}(y|X)$ and $P_{training}(X)=P_{testing}(X)$. We do not monitor data drift.
@@ II-B Semi-Supervised Concept Drift Detection for Machine Learning Systems
CDDs can be categorized into supervised, semi-supervised, and unsupervised methods. Supervised approaches require true labels for all data, which is often unrealistic. In contrast, unsupervised and semi-supervised approaches require true labels for either no data or only a few data points, making them more practical in real-world applications. Current unsupervised techniques detect concept drift only when it coincides with data drift, rendering them ineffective at detecting standalone concept drifts. The ability to detect concept drifts directly is desirable since there is no guarantee that data and concept drifts happen simultaneously.
Based on our industry needs, we believe semi-supervised approaches are suitable to detect concept drift accurately and efficiently in production. These approaches require minimal manual labeling and can target independent concept drifts. In a nutshell, our work focuses on proposing a better semi-supervised approach to detect concept drifts as we believe semi-supervised approaches are better aligned with our industry needs.
@@ III-A Concept Drift Detection
Supervised concept drift detection. Supervised CDDs use the true labels of all data points to detect concept drift. They either compare statistics of the classifier error rate between the current and reference windows [25, 26, 27, 28, 29, 30] or use the intersection of confidence intervals from multiple CDDs to detect concept drift [31, 32]. While these methods are fast and accurate, they require labels for all data points which is unrealistic in production given the velocity of data streams [33].
(Semi)-supervised concept drift detection. As a more production-friendly alternative, semi-supervised CDDs solve the problem by requiring only a subset of the data to be labeled. Existing methods can do this by using active learning [34, 35, 36], computational geometry [37], class imbalance [38], classifier confidence [39, 8], pseudo-error from a classifier ensemble [9], grid density sampling [40], ambiguity concerning the decision boundary and neighboring points [12], etc. to select only important data instances for labeling, reducing the labeling effort. However, the reliance on specific metrics to sample labels is vulnerable to data distribution and model architecture changes. Also, requiring the original model to follow certain ML architectures significantly limits adoption. We look to overcome these limitations.
Unsupervised concept drift detection. Unsupervised CDDs [10] detect drifts offline in batch [41, 42, 43, 44, 45, 46, 47, 48, 49, 50] or online by instances [51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. These methods detect concept drift by monitoring $P_{training}(X)\neq P_{testing}(X)$ under the assumption that changes in the data distribution and the posterior probability distributions often happen simultaneously [16]. Hence, unsupervised methods cannot detect concept drift when data drift is absent (Subsection II-A).  CDSeer aims to detect real concept drift even when the input data distribution stays the same.
@@ III-B Testing Machine Learning Systems
Artificial Intelligence for IT Operations (AIOps) AIOps leverage ML models for tasks such as anomaly detection and failure prediction [62, 63, 64, 3, 65, 66, 67, 5, 2, 68] using logs, measurements, and spatial information. These models are highly susceptible to data evolution, which impairs their performance and stability. Current mitigation strategies involve periodically retraining models [67, 68] or retraining based on alarms from supervised CDDs [6], both of which are costly in manual efforts due to the extensive labeling required. In this work, we explore CDSeer’s concept drift detection for AIOps to reduce cost.
Model Robustness and Relevance Testing. Model testing detects mismatches between the model and data [69, 70] and ensures trained models’ resistance to data noise [71, 72, 73]. However, these techniques test pre-trained models offline and more online techniques are needed [61]. CDSeer detect concept drift online to test the robustness and relevance of ML models. Thus  CDSeer can benefit the utilization and deployment of ML models for SE tasks. In this work, we applied CDSeer to an internal dataset for anomaly detection in router systems and the Google Cluster Trace dataset for job failure prediction.
@@ IV CDSeer: Learning-based Concept Drift Detection
In this section, we first summarize our approach at a high level. Then, we describe the framework step-by-step, with design choices and implementation details.
@@ IV-A An Overview of CDSeer
At a high level, CDSeer works entirely online as follows. As new data points arrive, the online model predicts labels in parallel with the inspector model predicting pseudo-labels. Concept drift is detected by a supervised CDD that monitors the distribution of classification error between the labels and pseudo-labels. The inspector model is trained as follows: the window of the latest data points is continuously clustered. From these clusters, a stratified sample is selected for labeling by experts and stored in memory. Via semi-supervised learning, the true labels in memory are propagated back to the window of the latest data points, where these newly generated labels and their features are used to update the inspector model. Finally, while this is not the main focus of our experiments, we follow convention and retrain the online model when concept drift is detected using true labeled data.
@@ 15    $(W,TrueLabels)\leftarrow$ spread labels from $M$ to $W$;
We utilize a supervised CDD to monitor the distribution of the $error$s rate and alert when it crosses a threshold, signaling the online and inspector models are trained using different concepts. For this purpose, the CDD can be chosen from well-known CDDs such as the Drift Detection Method [26], Page-Hinkley test (PHT) [27], etc. The CDD monitors the aggregated $error$ rate between the online and inspector models for increases surpassing a predefined threshold, which triggers a concept drift alarm between the online and the inspector models. Then, the framework records the corresponding instance and proceeds to the next data point.